OpenStudy (anonymous):

Identify the correct x intercept and y intercept of the equation 4x - 3y = 24. My answers is x intercept = (8, 0); y intercept = (0, 6) Am I right? :)

OpenStudy (e.cociuba):

To get the y-intercept, let x = 0: 4(0)-3y = 24, or -3y=24, or y = -8. Therefore, the y-intercept is the point (0, -8). To get the x-intercept, let y = 0: 4x - 3(0) = 24, or 4x = 24, or x = 6. Therefore, the x-intercept is the point (6, 0). This is wat i got!!

OpenStudy (evonhowell):

(-8,0) is x (0,6) is y I looked at my graph wrong.. :)

OpenStudy (anonymous):

so the answer is x intercept = (6, 0); y intercept = (0, -8)?

OpenStudy (e.cociuba):

yep :)

OpenStudy (anonymous):

Thank you so much! :)

OpenStudy (e.cociuba):

np:)

OpenStudy (evonhowell):